OpenStudy (anonymous):

Find three consecutive even numbers so that the sum of the first and third numbers is 22 less than three times the second number.

OpenStudy (larseighner):

let the smallest number be n. Then the three numbers are n, n+1, and n+2. n + (n+2) = 3(n+1) - 22 One equation in one unknown. Solve.

OpenStudy (larseighner):

Whoops EVEN numbers So make them n, n+2, n+4 n + (n+4) = 3(n+2) - 22

OpenStudy (anonymous):

thank you :)